### v2

# Adaptive Multi-Scale Integration Unlocks Robust Cell Annotation in Histopathology Images

Abstract
Identifying cell types and subtypes in routine histopathology is fundamental for understanding disease. Existing tile-based models captures nuclear detail but miss the broader tissue context that influences cell identity. Current human annotations are coarse-grained and uneven across studies, making fine-grained, subtype-level classificsation difficult. In this study, we build a marker-guided dataset from Xenium spatial transcriptomics with single-cell–resolution labels for more than two million cells across eight organs and 16 classes to overcome the lack of high-quality annotations. Leveraging this data resource, we introduce NuClass, a pathologist workflow–inspired framework for cell-wise multi-scale integration of nuclear morphology and microenvironmental context. It combines Path local, which focuses on nuclear morphology from $224^{2}$-pixel crops, and Path global, which models the surrounding $1024^{2}$-pixel neighborhood, through a learnable gating module to balance local and global details. An uncertainty-guided objective directs the global path to prioritize regions where the local path is uncertain, and we provide calibrated confidence estimates and Grad-CAM maps for interpretability. Evaluated on three fully held-out cohorts, NuClass achieves up to 96% F1 for its best-performing class, outperforming strong baselines. Our results demonstrate that multi-scale, uncertainty-aware fusion can bridge the gap between slide-level pathological foundation models and reliable, cell-level phenotype prediction.
Figure 1: NuClass: cell-wise multi-scale classification with gating. (Top) From whole-slide images (WSIs), we extract paired fields centered on each cell: a nucleus-scale crop ($224^{2}$ at 0.25 mpp) and a co-registered contextual field-of-view (FOV; default $1024^{2}$ at 0.25 mpp). Spatial assays (Xenium-like) provide gene-marker profiles and cell centroids, enabling scalable, subcellular annotations. (Stage 1, Path local) A UNI2-h morphology backbone processes the $224^{2}$ crop. A tissue-conditioned FiLM adaptor modulates features before a linear head outputs per-cell probabilities. (Stage 2, Path global) A DINOv3 ViT-L/16 encoder ingests the $1024^{2}$ contextual FOV; its representation is concatenated with a Path local morphology vector to predict complementary probabilities. (Stage 3, Fusion Gate) A lightweight cell-wise gate receives statistics from both distributions and compact feature projections, then fuses experts in probability space $\mathbf{p}_{\text{mix}}=(1-g)\mathbf{p}_{local}+g\mathbf{p}_{global}$. The gate learns which expert to trust per cell, not to average logits.
1 Introduction
Understanding cell types and their spatial interplay within tissues is central to studying human disease. While emerging technologies such as spatial transcriptomics enable cell type identification through marker genes, hematoxylin and eosin (H&E)-stained whole-slide images remain the most practical and widely adopted modality in routine clinical workflows, serving as the cornerstone for differential diagnosis and quantitative evaluation and guiding assessment of cancer stage, transplant outcomes, and treatment strategies.
In H&E images, nuclei segmentation [stardist] is well studied, but assigning detected nuclei to robust, generalizable cell identities—i.e., cell classification—remains challenging. This largely stems from (1) the lack of large-scale, cross-slice, high-quality, unified training datasets and (2) the difficulty of learning cell type–specific features under heterogeneous morphology across patients, organs, and scanners. Together, these issues reflect two core limitations: data scarcity and model design.
On the data side, recent studies show that a lack of high-quality, cell-level annotated training data, combined with stain/tissue heterogeneity and domain shifts, strongly constrains deep models for cell classification. Human annotations are often coarse and inconsistent and require substantial expert effort, limiting subcellular precision, scalability, and the diversity of labeled cell types.
On the modeling side, three challenges remain. (i) Current cell embedding models are largely unaware of a cell’s neighborhood, even though distinguishing, for example, cancer-associated from normal epithelial cells or macrophages from T cells often requires context beyond the local nuclear image. Multi-scale pipelines for gigapixel slides improve slide-level reasoning [chen2022hipt], but aggregation typically occurs at the bag or slide level, leaving the instance-wise balance between morphology and context underexplored. (ii) Models struggle with varying annotation granularities and taxonomies across tissues and datasets (e.g., the same cell may be labeled as “inflammatory cell,” “immune cell,” or “lymphocyte”). (iii) There is no large-scale pretrained foundation model across diverse diseases and organs to provide a strong encoder for downstream fine-tuning.
These spatial and semantic nuances across data, annotations, and models remain underexplored, creating an urgent need for a foundation model trained on millions of cells that can harmonize diverse annotations and enable deeper understanding of cell types and tissue organization.
To address data scarcity and model design jointly, we present a workflow from data curation to model development to advance understanding of cellular and nuclear features in H&E-stained images. On the data side, we build a scalable, marker-guided annotation pipeline on Xenium spatial assays that assigns fine-grained cell types with pixel-level coordinates to millions of cells, complementing and refining human supervision [xeniumtech]. The resulting labels are objective and reproducible across tissues and institutions and provide precise anchors for cross-scale modeling and evaluation (Fig. 1, top).
On the modeling side, we develop NuClass, an adaptive framework for cell-wise multi-scale integrated learning. NuClass uses two fixed fields of view: Path local focuses on nucleus-centered morphology using small crops, whereas Path global processes the co-registered larger neighborhood to capture tissue architecture and micro-environmental cues. A learnable gate decides, for each cell, how much to rely on local detail versus global context, and training encourages the two paths to specialize and complement rather than duplicate one another.
At scale, NuClass generalizes across organs and datasets for reliable cell-level phenotyping. Trained on more than 2M marker-labeled cells spanning 16 types across 8 organs and evaluated on three fully held-out cohorts (about 0.28M cells), it achieves up to 96% for its best-performing cell class and consistently matches or surpasses strong baselines while producing well-calibrated, interpretable predictions.
2 Related Work
Multi-organ multi-class H&E cell datasets. Recent dataset efforts have shifted from single-organ toward multi-organ collections, such as MoNuSAC [verma2021monusac] (4 tissues) and PanNuke [gamper2019pannuke] (19 organs/tissues). However, due to the scarcity of large, well-annotated cell datasets, most cell-classification models remain tailored to specific organs or narrow use-cases, as in HoVer-Net [graham2019hovernet] and CellViT [CellViT].
Foundation models for cell classification. Foundation models are reshaping computational pathology by learning general-purpose visual or vision–language representations from large, heterogeneous datasets [chen2024uni, xu2024gigapath, vorontsov2024virchow, lu2024conch, lu2024pathchat, ding2025titan, xiang2025musk]. These approaches transfer well on patch- and slide-level benchmarks, and vision–language systems extend to retrieval, question answering, and outcome prediction. Yet most progress still targets slide-level endpoints. In routine practice, a cell’s identity depends on both its nuclear morphology and the tissue neighborhood in which it resides, but existing systems seldom integrate these two sources of evidence at per-cell granularity. The recent VOLTA algorithm [nakhli2024volta] leverages over 800k cells across six cancer types to train a self-supervised cell representation model via contrastive learning, but it still relies on cell patches and spatial neighborhoods as unlabeled training signals and does not provide label-consistent, cross-organ cell types.
Per-cell representation learning at scale. The transition from patch-level to cell-level understanding marks a fundamental step toward morphological precision. Traditional architectures like HoVer-Net [graham2019hovernet] and CellViT [horst2024cellvit] jointly segment and classify cells, yet remain trained on a limited set of annotated organs. Graph- or structure-aware models [lou2024senc] extend this by modeling cell neighborhoods and propagating context among adjacent cells to capture micro-environmental cues. More recently, VOLTA [nakhli2024volta] proposed environment-aware contrastive learning, jointly encoding nuclear morphology and tissue surroundings to obtain self-supervised embeddings over 0.8M cells; however, these embeddings remain unlabeled and cannot be directly aligned across organs or histologic types. Beyond isolated cells, cross-organ datasets such as PanNuke [gamper2019pannuke, gamper2020pannuke_ext], Lizard [graham2021lizard], and NuCLS [amgad2022nucls] highlight the importance of organ diversity for robust per-cell modeling, yet per-cell recognition still struggles to unify morphology and context across domains. A scalable foundation model explicitly designed for per-cell representation—bridging nuclear texture, multi-scale tissue context, and cross-organ semantics—therefore remains an open problem in computational pathology.
3 Method
Given a cell–centric dataset with tissue annotations $\mathcal{D}=\{(\mathbf{x}^{224}_{i},\mathbf{x}^{ctx}_{i},t_{i},y_{i})\}_{i=1}^{N}$, $\mathbf{x}^{224}_{i}\!\in\!\mathbb{R}^{224\times 224\times 3}$ denotes a local $224^{2}$ patch, $\mathbf{x}^{ctx}_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$ (H=W=1024H{=}W{=}1024 by default) a wider contextual view, $t_{i}$ the tissue type (e.g., liver/lung), and $y_{i}\in\{1,\dots,C\}$ the target cell class. We introduce two complementary branches: Path local focuses on cell morphology with tissue-aware adaptation, while Path global injects context to cover Path local’s blind spots. For fusion, we freeze the two branches and train a fusion gate to combine predictions in probability space, yielding stable gains without modifying either backbone.
Notation. For a $C$-class softmax output we write $\mathbf{p}\in\Delta^{C-1}$ to mean that $\mathbf{p}=(p_{1},\dots,p_{C})^{\top}\in\mathbb{R}^{C}$ lies on the probability simplex $\Delta^{C-1}=\{\,\mathbf{u}\in\mathbb{R}^{C}:\ u_{c}\geq 0,\ \sum_{c}u_{c}=1\,\}$. We denote by $p_{\max}=\max_{c}p_{c}$ the largest class probability (“confidence”), by $\mathrm{H}(\mathbf{p})=-\sum_{c=1}^{C}p_{c}\log p_{c}$ the Shannon entropy [shannon1948], and by $\mathrm{margin}(\mathbf{p})=p_{(1)}-p_{(2)}$ the gap between the largest and second–largest coordinates (a simple confidence margin). For an input example $x_{i}$, $\mathbf{p}_{\text{local}}$ and $\mathbf{p}_{\text{global}}$ are the softmax probability vectors predicted by Path local and Path global, respectively; $\mathbf{p}_{\text{mix}}$ denotes the final fused distribution (a convex combination used at test time).
3.1 Path local: Tissue-aware Cell Morphology Classifier
Motivation and task. In high-throughput pathology, morphology is the most direct signal, path local aims to learn robust, transferable morphology from $224^{2}$ patches by adapting to tissue domains.
Architecture. We adopt a UNI2-h ViT backbone $\mathcal{F}_{uni}$ [chen2024uni, dosovitskiy2021vit] to extract a global feature $\mathbf{h}_{i}=\mathcal{F}_{uni}(\mathbf{x}^{224}_{i})\in\mathbb{R}^{1536}$. LayerNorm [ba2016layernorm] and FiLM [perez2018film] (conditioned on tissue) stylize the feature:
$\mathbf{e}_{t}$ $=\mathrm{Emb}(t_{i})\in\mathbb{R}^{d_{t}},$ (1)
$(\boldsymbol{\gamma}_{i},\boldsymbol{\beta}_{i})$ $=\phi(\mathbf{e}_{t})\in\mathbb{R}^{1536}\times\mathbb{R}^{1536},$ (2)
$\tilde{\mathbf{h}}_{i}$ $=\mathrm{LN}(\mathbf{h}_{i})\odot\bigl(1+\boldsymbol{\gamma}_{i}\bigr)+\boldsymbol{\beta}_{i},$ (3)
where $\phi(\cdot)$ is a 2-layer MLP and $\odot$ denotes elementwise product. A 2-layer MLP head outputs logits and probabilities: $\mathbf{z}_{i}^{\text{local}}=\mathrm{Head}_{\text{local}}(\tilde{\mathbf{h}}_{i})\in\mathbb{R}^{C},\ \mathbf{p}_{\text{local}}=\mathrm{softmax}(\mathbf{z}_{i}^{\text{local}})$.
Loss and schedule. We use class-balanced cross-entropy with label smoothing [cui2019classbalanced, muller2019label_smoothing]:
$\mathcal{L}_{\text{local}}=-w_{y_{i}}\!\!\sum_{c=1}^{C}\!q_{i,c}\log p_{\text{local},c},\quad q_{i,c}=\begin{cases}1-\epsilon,&c=y_{i},\\ \frac{\epsilon}{C-1},&c\neq y_{i},\end{cases}$ (4)
with class weights from the Effective Number: $w_{c}\!\propto\!\frac{1-\beta}{1-\beta^{n_{c}}}$ ($n_{c}$: samples of class $c$). We adopt a probe warm-up followed by full fine-tuning [kolesnikov2020bit], and apply layer-wise LR decay to the ViT for stability [bao2022beit, touvron2022deit3].
Path local emphasizes cell-intrinsic morphology. Yet $224^{2}$ local windows can be insufficient under complex micro-environments or cell–cell interactions, motivating Path global to supply context.
3.2 Path global: Context-aware Complementary Branch
Motivation and task. Many cell types become separable only when neighborhood cues (e.g., glands, vessels, immune infiltration) are visible; a $224^{2}$ crop alone is often insufficient to tell them apart. Path global injects wide-field context while preserving Path local’s morphology competence, and—crucially—learns where Path local is weak instead of relearning what Path local already solves.
Local-initialized backbone. Path global reuses the UNI2-h backbone from Path local on the $224^{2}$ crop. We initialize UNI, the tissue-aware FiLM, and the local-style head from Path local’s checkpoint. For the first $E_{\text{freeze}}$ epochs these modules are frozen, preventing early drift and ensuring Path global starts from Path local’s morphology.
Architecture (with tissue FiLM in context). Given a crop $\mathbf{x}^{224}_{i}$ and a contextual view $\mathbf{x}^{ctx}_{i}\!\in\!\mathbb{R}^{H\times W\times 3}$, we build a two-stream encoder:
$\mathbf{u}_{i}$ $=\mathcal{F}_{uni}(\mathbf{x}^{224}_{i})\in\mathbb{R}^{1536},$ (5)
[…]
apply tissue-conditioned FiLM to the context stream,
$(\boldsymbol{\gamma}^{\text{ctx}}_{i},\boldsymbol{\beta}^{\text{ctx}}_{i})$ $=\phi_{\text{ctx}}(\mathbf{e}_{t}),$ (7)
$\tilde{\mathbf{v}}_{i}$ $=\mathrm{LN}(\mathbf{v}_{i})\odot\bigl(1+\boldsymbol{\gamma}^{\text{ctx}}_{i}\bigr)+\boldsymbol{\beta}^{\text{ctx}}_{i},$ (8)
then project and concatenate:
$\mathbf{u}^{\prime}_{i}$ $=\mathrm{LN}(\mathbf{u}_{i})\mathbf{W}_{u}\in\mathbb{R}^{512},$
$\mathbf{v}^{\prime}_{i}$ $=\mathrm{LN}(\tilde{\mathbf{v}}_{i})\mathbf{W}_{v}\in\mathbb{R}^{512},$
$\mathbf{s}_{i}$ $=[\,\mathbf{u}^{\prime}_{i};\,\mathbf{v}^{\prime}_{i}\,]\in\mathbb{R}^{1024},$ (9)
$\mathbf{z}^{\text{global}}_{i}$ $=\mathrm{Head}_{\text{global}}(\mathbf{s}_{i}),\qquad\mathbf{p}^{\text{global}}_{i}=\mathrm{softmax}(\mathbf{z}^{\text{global}}_{i}).$ (10)
Why this fusion is efficient. (i) Both streams are LayerNorm’ed and projected to a common 512-d space, so their scales are comparable; (ii) the 1024-d concatenation lets $\mathrm{Head}_{\text{global}}$ learn the mixture in feature space, which is expressive yet compute-efficient; (iii) empirically this “LN+linear-to-512, then concat” is low-cost and consistently effective.
Context resolution and resampling. Unless specified otherwise, the context field-of-view (FOV) is a $1024^{2}$ square centered at the target cell at native slide resolution. For compute efficiency we may resample the same FOV to $r_{\text{ctx}}\!\in\!\{512,768,1024\}$ (default $r_{\text{ctx}}{=}1024$) before DINO. Resampling preserves the spatial FOV—no cropping/padding—only the sampling density changes. If an integer ratio exists (e.g., $1024{\rightarrow}512$) we use stride decimation; otherwise bilinear interpolation. With patch size 16, DINO processes $S=(r_{\text{ctx}}/16)^{2}$ tokens and attention scales as $\mathcal{O}(S^{2})$ [vaswani2017attention]; we therefore employ micro-batched forward (e.g., 16–64) to bound memory and compilation latency without sacrificing throughput. For cells near slide boundaries, out-of-range regions are mirrored or reflected so that the effective FOV remains $1024^{2}$ without introducing artificial padding or empty borders.
Local-aware weighting (definition). Let $\mathbf{z}^{\text{local}}_{i}$ be Path local head logits and $\mathbf{p}^{\text{local}}_{i}=\mathrm{softmax}(\mathbf{z}^{\text{local}}_{i})$. We up-weight samples for which Path local is uncertain by
$w_{i}\;=\;\bigl(1-p^{\text{local}}_{i}(y_{i})\bigr)^{\gamma}$ (11)
so $w_{i}$ increases when Path local assigns a lower probability to the ground-truth class; this weighting mirrors the focal reweighting pattern [lin2017focal].
Complementarity supervision (losses). Our default objective for Path global is
$\mathcal{L}_{\text{main}}$ $=\mathbb{E}_{i}\Big[w_{i}\cdot\ell\big(\mathbf{z}^{\text{global}}_{i},y_{i}\big)\Big],$ (12)
$\mathcal{L}_{\text{stable}}$ $=\mathbb{E}_{i}\big[\ell(\mathbf{z}^{\text{global}}_{i},y_{i})\big],$ (13)
$\mathcal{L}_{\text{PathGlobal}}$ $=\mathcal{L}_{\text{main}}\;+\;\lambda_{\text{stable}}\,\mathcal{L}_{\text{stable}},$ (14)
[…]
Regularizers: “use context when helpful”. We apply two light regularizers to avoid collapse to the crop stream: (i) Context dropout ($\mathbf{v}^{\prime}_{i}\!\leftarrow\!\mathbf{0}$ with prob. $\rho$), encouraging Path global to use context only when present; (ii) Context warm-up, concatenating $\mathbf{u}^{\prime}_{i};\,\alpha\,\mathbf{v}^{\prime}_{i}\,]$ with $\alpha$ ramped from $\alpha_{0}$ to $1$ over the first $E_{\alpha}$ epochs. These follow dropout/stochastic depth and curriculum learning [srivastava2014dropout, huang2016stochasticdepth, bengio2009curriculum].
Why it yields complementarity. Eq. (11) shifts gradient mass to the Path local-hard region (small $p^{\text{local}}_{i}(y_{i})$), while Eq. (13) keeps Path global calibrated on easy cases. Starting from a Path local–initialized (and briefly frozen) morphology backbone avoids early drift; warm-up/dropout ensure Path global exploits context only when it helps. Together, Path global learns contextual evidence Path local lacks rather than duplicating morphology decisions.
3.3 Gated Fusion of Path local and Path global
Motivation and formulation. Path local and Path global exploit different signals—fine nuclear morphology vs. wide-field context—and their logits can reside on different scales with distinct uncertainty profiles. Mixing logits is brittle. We therefore fuse in probability space via a scalar, per-cell gate $g_{i}\!\in\![0,1]$ (mixture-of-experts):
$\mathbf{p}_{\text{mix},i}\;=\;(1-g_{i})\,\mathbf{p}_{\text{local},i}\;+\;g_{i}\,\mathbf{p}_{\text{global},i},$ (15)
which is numerically stable and cannot overshoot either path.
Gate inputs and architecture (GT-free). The gate reads two compact feature projections and GT-free distributional statistics. Let $\mathbf{f}^{\text{local}}_{i}\!\in\!\mathbb{R}^{1536}$ be Path local’s pre-head feature and $\mathbf{s}^{\text{global}}_{i}\!\in\!\mathbb{R}^{1024}$ be Path global’s fused feature (UNI patch + DINO context). We form
$\tilde{\mathbf{f}}^{\text{local}}_{i}$ $=\mathrm{LN}(\mathbf{f}^{\text{local}}_{i})\mathbf{W}^{\text{local}}\in\mathbb{R}^{d_{\text{local}}},$ (16)
$\tilde{\mathbf{s}}^{\text{global}}_{i}$ $=\mathrm{LN}(\mathbf{s}^{\text{global}}_{i})\mathbf{W}^{\text{global}}\in\mathbb{R}^{d_{\text{global}}}.$
We then assemble a statistic vector
$r_{i}$ $=[\,p^{\text{local}}_{\max},\,p^{\text{global}}_{\max},\,\bar{\mathrm{H}}(\mathbf{p}_{\text{local}}),\,\bar{\mathrm{H}}(\mathbf{p}_{\text{global}}),$ (17)
$\quad\mathrm{margin}(\mathbf{p}_{\text{local}}),\,\mathrm{margin}(\mathbf{p}_{\text{global}}),\,\rho_{i},\,|\rho_{i}|\,],$
where $p^{\cdot}_{\max}=\max_{c}p^{\cdot}_{c}$ is top‑1 confidence, $\bar{\mathrm{H}}(\cdot)$ is entropy normalized by $\log C$, $\mathrm{margin}(\cdot)=p_{(1)}-p_{(2)}$, and
$\rho_{i}\;\triangleq\;p^{\text{global}}_{\max}-p^{\text{local}}_{\max}$
captures GT-free confidence disagreement between the two paths. A two-layer MLP $\psi(\cdot)$ outputs
$g_{i}=\sigma\!\big(\psi\big([\tilde{\mathbf{f}}^{\text{local}}_{i};\tilde{\mathbf{s}}^{\text{global}}_{i};r_{i}]\big)\big).$ (18)
Supervision: “choose whom to trust” (training-only use of GT). To supervise the gate, we define predicted labels $\hat{y}_{\text{local}}=\arg\max_{c}p^{\text{local}}_{i,c},\;\hat{y}_{\text{global}}=\arg\max_{c}p^{\text{global}}_{i,c}$ and a soft preference using the ground-truth class:
$\Delta^{(y)}_{i}$ $\;\triangleq\;p^{\text{global}}_{i}(y_{i})-p^{\text{local}}_{i}(y_{i}),$
[…]
Crucially, $\Delta^{(y)}_{i}$ is used only for supervision and loss weighting during training; it is never part of the gate input $r_{i}$ and is not needed at test time.
Conflict-aware weighting (training-only). We emphasize samples on which the two paths disagree, and modulate by their confidence gap on the ground-truth class:
$w_{i}\;=\;1+\lambda_{\text{conf}}\cdot\mathbf{1}\!\big[(\hat{y}_{\text{local}}{=}y_{i})\oplus(\hat{y}_{\text{global}}{=}y_{i})\big]+\lambda_{\Delta}\,|\Delta^{(y)}_{i}|^{\gamma_{\Delta}}.$ (19)
Objective. The training loss combines fused NLL, gate BCE, a consistency alignment, and an entropy penalty:
$\mathcal{L}_{\text{mix}}$ $=\mathbb{E}_{i}\!\big[\,w_{i}\cdot\big(-\log\mathbf{p}_{\text{mix},i}(y_{i})\big)\,\big],$ (20)
$\mathcal{L}_{\text{gate}}$ $=\mathbb{E}_{i}\!\big[\,w_{i}\cdot\mathrm{BCE}\big(g_{i},\tilde{g}_{i};\,\mathrm{pos\_w}\big)\,\big],$ (21)
$\mathcal{L}_{\text{align}}$ $=\mathbb{E}_{i}\!\big[\,w_{i}\cdot\mathrm{KL}\big(\mathbf{p}^{\star}_{i}\,\|\,\mathbf{p}_{\text{mix},i}\big)\,\big],$ (22)
$\mathcal{L}_{\text{ent}}$ $=\mathbb{E}_{i}\!\big[\,\lambda_{\text{ent}}\cdot|\Delta^{(y)}_{i}|\cdot\mathrm{H}_{\text{Bern}}(g_{i})\,\big],$ (23)
where $\mathbf{p}^{\star}_{i}=\mathbf{p}_{\text{global},i}$ if $\tilde{g}_{i}>0.5$ and $\mathbf{p}^{\star}_{i}=\mathbf{p}_{\text{local},i}$ otherwise, and $\mathrm{pos\_w}$ is an adaptive positive weight from the batch. The full objective is
$\mathcal{L}_{\text{gate-fuse}}\;=\;\mathcal{L}_{\text{mix}}+\mathcal{L}_{\text{gate}}+\lambda_{\text{align}}\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{ent}}+\lambda_{\text{aux}}\,\mathcal{L}_{\text{aux}},$ (24)
with a small auxiliary term $\mathcal{L}_{\text{aux}}$ added only when unfreezing heads/backbones.
Schedule and deployment (safe gate; GT-free). We first train the gate with Path local/Path global frozen (gate-only warm-up), then optionally unfreeze lightweight modules with small LRs. At test time we use Eq. (15). For deployment we provide a safe gate calibrated on validation data: compute per-path reliability by predicted class, form a baseline chooser by $p_{\max}\!\times$ reliability, and accept the gate’s choice only when
$g_{i}\;>\;\tau\quad\text{and}\quad|\rho_{i}|\;>\;\gamma,$
with $\rho_{i}=p^{\text{global}}_{\max}-p^{\text{local}}_{\max}$. This conservative policy is GT-free and empirically avoids regressions vs. the best single path while often improving accuracy.
Figure 2: Organ and cohort wise composition. Donut plots of the training set (8 organs) and the hold-out testing cohorts (pancreas, ovary, lung). The figure illustrates the distribution shifts between training and evaluation cohorts.
Metric FiLM Concat $\Delta$(F–C)
ACC 0.8493 0.4845 +0.3648
F1-macro 0.5929 0.4076 +0.1853
Silhouette $\uparrow$ 0.1480 0.0182 +0.1298
CH $\uparrow$ 236.8 200.6 +36.2
DB $\downarrow$ 4.26 5.56 –1.30
ECE $\downarrow$ 0.061 0.213 –0.152
Table 1: FiLM improves feature separability and probability calibration without changing the training pipeline. Silhouette/CH/DB follow the classic definitions [rousseeuw1987silhouette, calinski1974ch, davies1979db]; ECE follows [guo2017calibration].
4 Experiments
4.1 Experimental Setup
Datasets and splits. All methods are trained on the same multi-organ dataset covering breast, colon, heart, kidney, liver, lung, ovary, and pancreas, and evaluated on three disjoint hold-out Xenium cohorts (lung, ovary, and pancreas). Each image originates from a distinct patient, so there is no patient-level overlap between training and testing cohorts, and no images or annotations from any hold-out cohort are seen during training. To substantiate the gene-marker-based labeling scheme, Appendix lists all cell types together with their top $10$ marker genes, providing an additional justification for our annotation strategy. Figure 2 summarizes the organ and cohort composition.
[…]
$\mathbf{p}_{\text{eval}}\;=\;\mathbf{M}^{\top}\mathbf{p}_{\text{train}},$ (25)
where each column of $\mathbf{M}$ sums probabilities of training classes that map to the same evaluation class. For each cohort, the evaluation label set is restricted to the intersection between the canonical taxonomy and the cell types present in that cohort, so that all methods are evaluated on the same set of labels.
Information parity & implementation. To equalize metadata usage, all fine-tuned models (ours and baselines) attach a learned 32-d organ embedding to the visual feature and train it jointly with the classifier, [^1] Unless noted otherwise, we concatenate the organ embedding to the visual feature. while DINO baselines use frozen backbones with a linear probe and therefore do not use organ embeddings, preserving the standard LP protocol. Fine-tuned runs share the same optimizer (AdamW), cosine learning-rate decay with warmup, early stopping on validation macro–F1, PyTorch 2.x with BF16 (TF32 on Ampere+), and identical dataloaders, precision settings, batch sizes, and training schedules per resolution [loshchilov2019adamw, loshchilov2017sgdr, goyal2017imagenet, paszke2019pytorch].
4.2 Comparison with SOTA Models
Baselines and training setup.
All baselines are instantiated in the unified label space and data pipeline described above. For MUSK, LOKI, and PLIP we start from the official checkpoints and recommended training recipes from their original papers and public code, adjusting only image resolution, batch size, and training epochs to match the available GPU memory and keep the total number of optimization steps within the same order of magnitude as our method. The resulting configurations and GPU usage are summarized in Appendix.
Our model. We evaluate all three stages of our framework on the same hold-out cohorts: (1) a Path local expert specialized for small-scale, densely structured morphology at $224^{2}$ resolution; (2) a Path global expert capturing large-field contextual patterns at $1024^{2}$; and (3) a learned Fusion Gate that adaptively integrates the two paths by weighting logits per input. Hyperparameters and training details for these three components are provided in Appendix.
[…]
Table 2: Per-class F1 on hold-out datasets. Legend: DINO-B/L/S+/16 = DINOv3 ViT-B/L/S(16+); Path local and Path global are our single-path experts; Fusion Gate is our learned gating fusion. Best/second/third are marked as  / / .
SOTA models Ours
Frozen backbone + LP Fully fine-tuned
Organ Cell type DINO-B/16 DINO-L/16 DINO-S+/16 LOKI PLIP MUSK Path local Path global Fusion Gate
Lung Epithelial Malignant 84.62 83.40 48.87 86.34 65.62 81.35 93.90 95.23 96.42+1.18
Endothelial 50.18 55.66 36.13 56.04 51.08 73.17 82.49 83.53 83.76+0.24
T Cell 9.72 9.28 11.99 25.49 21.73 28.18 13.23 43.89 54.06+10.17
CAFs 35.48 24.29 9.55 25.67 19.41 28.91 47.55+2.69 43.74 44.86
Ovary Epithelial Malignant 90.33 90.64 82.87 64.28 87.36 83.57 94.14 93.86 94.17+0.03
Macrophage 9.20 10.33 17.44 17.69 16.70 30.17 20.78 34.78 35.15+0.37
Endothelial 22.00 20.71 7.40 28.87 35.57 53.25 75.20 73.44 77.96+2.76
Pancreas Endothelial 9.68 6.63 8.33 15.28 19.67 22.49 49.38+0.66 48.44 48.72
T Cell 9.53 9.59 5.41 7.26 9.01 7.51 20.11 21.36 22.54+1.18
Fibroblast 46.53 45.37 60.86 57.48 13.20 65.14 71.58+2.94 62.87 68.65
Endocrine 0.00 0.00 8.92 0.00 1.24 20.26 87.39+6.54 80.85 68.90
Ductal 4.29 0.00 0.00 0.00 1.66 4.12 39.24+4.00 31.33 35.24
Acinar 0.00 0.00 16.46 0.00 1.01 18.97 44.30 42.65 44.60+0.30
Results summary and takeaways.
Table 2 reports per-class F1 across all hold-out cohorts. Overall, our Path local and Path global experts remain strong and complementary, and the learned Fusion Gate typically matches or exceeds the better single path. For example, on lung, Fusion Gate improves malignant epithelium to 96.42 F1 and raises T cells from 43.89 (global) to 54.06 (+10.17+10.17). On ovary, it increases endothelial from 75.20/73.44 (local/global) to 77.96. These gains reflect the benefit of combining fine-grained morphology with wide-field tissue context.
At the same time, fusion is not uniformly superior to the best single path. In pancreas, several morphology-dominant classes are better handled by Path local: endocrine (87.39 local vs. 68.90 fusion), fibroblast (71.58 vs. 68.65), and ductal (39.24 vs. 35.24); in lung CAFs, local (47.55) also exceeds fusion (44.86). We therefore (i) report per-cohort macro/micro F1 and confidence intervals in the Appendix, and (ii) deploy a calibrated safe gate that defers to the reliability-weighted single path unless the gate is confident and the two paths strongly disagree. Empirically this policy is no worse than the best single expert while often preserving the fusion gains on context-sensitive classes.
Takeaways. (1) Classes requiring architectural or microenvironmental cues (e.g., lymphocytes, ductal/endocrine, endothelial variants) benefit the most from global context and gated fusion. (2) For classes where nuclear morphology is decisive, the gate learns to favor Path local; safe gating further guards against over-fusion. (3) Under strong class imbalance and distribution shift (e.g., pancreas), multi-scale specialization with adaptive fusion remains robust on average, while acknowledging class-wise variability.
4.3 Ablation Studies
Tissue-aware FiLM in Path local. Replacing simple concatenation with a FiLM modulation between backbone and head improves separability and calibration [perez2018film]. Given feature $\mathbf{f}\!\in\!\mathbb{R}^{1536}$ and tissue embedding $\mathbf{e}_{t}\!\in\!\mathbb{R}^{64}$, FiLM predicts $(\gamma,\beta)=g_{\theta}(\mathbf{e}_{t})$ and applies $\tilde{\mathbf{f}}=(1+\gamma)\odot\mathbf{f}+\beta$. On a held-out lung subset, FiLM yields higher accuracy and macro–F1 and lower ECE (Table 1), with tighter clusters and a near-diagonal reliability curve (Fig. 3).
Figure 3: Lung subset: reliability and feature geometry. FiLM (left) vs. Concat (right). FiLM reduces ECE [guo2017calibration] (0.061 vs. 0.213) and produces more compact, tissue-aligned clusters.
Complementarity of scales (local/global). Evaluating the best checkpoints of Path local and Path global on eight validation sets (16 classes total) shows clear disagreement structure: Path global dominates in 11/16 classes that require broader context (e.g., cardiomyocyte, ductal, endocrine, T cell), while Path local leads in 4/16 morphology-dense classes (e.g., B cell, endothelial, smooth muscle, malignant epithelium); acinar is roughly balanced. An oracle that picks the correct expert per sample yields a positive upper bound, explaining the strong gains of Fusion Gate (Fig. 4).
Figure 4: Within-class shares of local-only (Path local correct, Path global wrong) vs global-only (Path global correct, Path local wrong). The asymmetry indicates complementary failure modes, motivating gated fusion.
Fusion-gate loss design. The fusion gate is trained with the composite loss $L=L_{\text{mix}}+L_{\text{gate}}+L_{\text{align}}+L_{\text{ent}},$ where $L_{\text{mix}}$ applies cross-entropy to the fused prediction, $L_{\text{gate}}$ supervises a soft target that favors the more reliable expert, $L_{\text{align}}$ encourages the fused distribution to stay close to the selected expert, and $L_{\text{ent}}$ regularizes the gate’s entropy. We evaluate the contribution of each component through a drop-one ablation on the eight-organ validation set, using precomputed features and training only the gate and classification heads. Table 3 reports the detailed results.
$L_{\text{mix}}$ $L_{\text{gate}}$ $L_{\text{align}}$ $L_{\text{ent}}$ Gate Safe ECE (Gate / Safe)
Acc F1 Acc F1 Gate Safe
0.8641 0.7843 0.8645 0.7617 0.0304 0.0603
$\times$ 0.0001 0.0000 0.0001 0.0000 0.6400 0.6515
$\times$ 0.8663 0.7880 0.8647 0.7794 0.0221 0.0333
$\times$ 0.8662 0.7867 0.8552 0.7079 0.0298 0.0951
$\times$ 0.8635 0.7831 0.8640 0.7572 0.0285 0.0606
[…]
4.4 Visualization
To examine where different models focus, we visualize token-level Grad-CAM overlays from the last transformer block on representative $224^{2}$ ovary patches (Fig. 5[selvaraju2017gradcam]. Across all examples, Path local produces compact, high-intensity responses that align with histologically meaningful structures, such as clusters of nuclei, epithelial–stromal interfaces, and sharp tissue borders. In contrast, LOKI, MUSK, and PLIP often yield more diffuse or fragmented activations that primarily follow coarse color/texture patterns and frequently extend into stromal or background regions with limited diagnostic value. This tighter spatial correspondence between Path local attention and nuclear morphology/boundaries is consistent with its stronger per-class F1 performance. More examples can be found at Appendix.
Figure 5: Token-level Grad-CAM on ovary patches. Columns: input, Path local, LOKI, MUSK, PLIP. Overlays (last block, $\alpha{=}0.45$) on $224^{2}$ patches. Compared with the baselines, Path local concentrates its responses on clusters of nuclei and tissue interfaces, whereas LOKI, MUSK, and PLIP often highlight broad texture patterns or even background regions, indicating weaker alignment with diagnostic structures.
5 Conclusion and Discussion
We presented NuClass, a cell-wise multi-scale framework that treats nuclear morphology and tissue context as independent experts and fuses them through a GT-free, probability-space gate. This design cleanly separates expert specialization from fusion, offers stable training via freeze–then–optional-unfreeze scheduling, and enables a conservative, reliability-calibrated deployment rule. Across three fully held-out cohorts, NuClass consistently matches or exceeds strong frozen and fine-tuned baselines, achieving up to 96% F1 for its best-performing class and maintaining robust generalization under distribution shift. Our experiments reveal three central findings:
  • Scale complementarity. Local morphology and global context excel on different phenotypes; the oracle gains (Fig. 4) indicate substantial untapped complementarity.
  • Adaptive but selective fusion. Fusion improves context-sensitive classes, whereas morphology-dominant classes remain better served by the local path, underscoring the need for per-class evaluation.
  • Safe deployment. A reliability-aware gating policy preserves nearly all fusion benefits while preventing regressions below the strongest single expert.
NuClass is directly usable as a plug-in module for existing cell-level pipelines and provides a principled mechanism for deferral when confidence is low. Nonetheless, limitations remain: marker-guided labels inherit Xenium biases; context modeling increases computational cost; and the dataset, while multi-organ, still lacks broader staining, scanner, and population diversity.
Future work will expand NuClass toward pan-cancer coverage, refine taxonomies to finer subtypes and cellular states, extend to end-to-end H&E-only pipelines with automatic nucleus detection, and explore adaptive context fields, efficient distillation, and stronger uncertainty calibration for clinical deployment. We also aim to incorporate active learning and uncertainty-guided sampling to more efficiently expand the training corpus with high-value annotations and reduce reliance on large-scale marker-based labeling.
Adaptive Multi-Scale Integration Unlocks Robust Cell Annotation
in Histopathology Images
Supplementary Material
Overview of the Supplementary Material
  • Baselines: provenance and rationale (Sec. A; Tab. A1):
    we describe how DINOv3 variants (frozen linear probe), PLIP, LOKI, and MUSK are instantiated under a unified protocol, including backbone choices, training protocol (frozen features vs. fine-tuning), and the rationale for each baseline family.
  • Full training and inference hyperparameters (Sec. B; Tabs. A2A4):
    we provide complete optimization settings for the path-local expert, path-global expert, and Fusion Gate, covering optimizers, learning rates, layer-wise decay, warm-up schedules, batch size / precision, data augmentation, tissue conditioning (FiLM), loss formulations, and early stopping. We also describe the organ/tissue prior and the “unknown organ” ablation switch used at inference time.
  • Safe Gate calibration and deployment (Sec. B, following Tabs. A2A4):
    we detail the counting of experts and the Safe Gate training configuration, and explain the validation-only procedure used to select conservative operating points for the gating policy at test time.
  • Compute budget and complexity (Tab. A5):
    we report parameter counts (trainable vs. non-trainable), total parameter size, peak GPU memory usage, wall-clock training time, and data scale for the local, global, and gate variants, together with the explicit counting policy that explains why Gate has $\sim 1.67$B non-trainable parameters.
  • Dataset composition, cohort splits, and marker genes (Sec. C; Tabs. A6A8):
    we summarize per-cohort cell counts, the number of unique cell types, and the train vs. hold-out composition after label rebuilding, and we list representative class-level marker genes used for molecular cell-type annotation.
[…]
    we formalize the canonical projection $p_{\text{eval}}=M^{\top}p_{\text{train}}$, provide a minimal numeric example (including tensor shapes), and give explicit cohort-wise binary projection matrices $M_{\text{lung}}$, $M_{\text{ovary}}$, and $M_{\text{pancreas}}$ that map the canonical training taxonomy to each evaluation taxonomy.
  • Cohort-wise Macro/Micro-F1, accuracy, and local-aware gating (Sec. E; Sec. E.1; Tab. A11):
    we report macro-F1 and micro-F1 (accuracy) with 95% confidence intervals from non-parametric bootstrap (B=1000B{=}1000) for each held-out cohort and overall, and analyze how the local-aware gating signal affects the global path in terms of accuracy, macro-F1, calibration (ECE, Brier score), AUROC, and performance on low-confidence / high-entropy subsets.
  • Additional visualizations (Sec. Additional visualizations; Figs. A1A2):
    we provide additional Grad-CAM visualizations for the path-local expert and for LOKI, MUSK, and PLIP on representative examples, illustrating qualitative differences in attention patterns across methods.
Appendix A Baselines: provenance and rationale
Model Type Backbone Training protocol Rationale DINOv3-B/16 SSL ViT (frozen) ViT-B/16 Frozen features + linear probe Canonical LP baseline DINOv3-L/16 SSL ViT (frozen) ViT-L/16 Frozen features + linear probe Larger-scale LP variant DINOv3-S+/16 SSL ViT (frozen) ViT-S+/16 Frozen features + linear probe Smaller LP variant PLIP VLP (fine-tuned) ViT-L/14 Vision tower fine-tuning Pathology VLP transfer LOKI VLP (fine-tuned) ViT-L/14-like Vision tower fine-tuning Visual–omics VLP baseline MUSK Vision-only (FT) p16 transformer End-to-end fine-tuning Strong non-VLP foundation model
Table A1: Snapshot of baseline families and protocols used in our experiments. All fine-tuned baselines employ the same 32-d organ embedding, data pipeline, and optimization setup for parity; DINOv3 variants are evaluated with frozen backbones and a learned linear probe.
Appendix B Full training and inference hyperparameters
Backbone UNI2-h (ViT-L/14-like, $224^{2}$, embed=1536) Optimizer / Weight Decay AdamW / $2{\times}10^{-5}$ LR (UNI / Head) $3{\times}10^{-6}$ / $3{\times}10^{-4}$ Layer-wise LR decay $\gamma$ $0.85$ (earlier layers $\to$ lower LR) Scheduler Linear warm-up to 1.0 (steps=$1500$); no decay Epochs / Freeze $50$ / freeze backbone for the first epoch Batch (effective) / Precision $512{\times}2{=}1024$ / BF16-mixed Data augmentation H/V flip, $\pm 15^{\circ}$ rotation, color jitter, affine, normalization Tissue conditioning FiLM (embedding dim $64$) Loss Class-balanced CE (Effective Number, $\beta{=}0.9999$) + label smoothing $\varepsilon{=}0.1$ Early stopping / Validation cadence patience $4$ / every $0.25$ epoch
Table A2: Path-local hyperparameters.
Backbones UNI2-h ($224^{2}$) + DINOv3-L/16 (context) Optimizer / Weight Decay AdamW / $2{\times}10^{-5}$ LR (UNI / Head / DINO) $2{\times}10^{-6}$ / $3{\times}10^{-4}$ / $1{\times}10^{-6}$ (DINO frozen by default) Layer-wise LR decay $\gamma$ $0.85$ (UNI) Scheduler Linear warm-up (steps=$2000$); no decay Epochs / Precision $40$ / BF16-mixed Batch (effective) $256{\times}2{=}512$ Context size / micro-batch $1024$ / $16$ (DINO forward chunking) Fusion $\alpha$ warm-up $\alpha{:}\;0.25\to 1.0$ (first $2$ epochs) Context dropout $0.10$ Freeze policy UNI unfrozen; DINO frozen (optional unfreeze) Loss $w=(1-p_{A}(y))^{\gamma},\ \gamma{=}2.0$; $\mathcal{L}=\mathcal{L}_{\text{main}}+0.3\,\mathcal{L}_{B}$
Table A3: Path-global hyperparameters.
[…]
Table A4: Fusion Gate hyperparameters.
Organ/tissue prior and ablation note.
In routine surgical pathology, organ/tissue identity is recorded at accession and remains available during H&E processing; we therefore treat it as known metadata and encode it via a 32-d embedding for all fine-tuned models (“information parity”). To simulate “unknown organ” without retraining, our released inference code provides a switch that replaces the organ embedding with a learned [UNK] token (or a zero-vector through FiLM), allowing users to assess sensitivity to this prior.
Table A5: Compute and memory budget of the local, global, and gate variants. Parameter counts follow PyTorch Lightning summaries; memory numbers are peak usage on a single H200 (141 GB).
Setup Trainable params Non-trainable params Total params Param size local (baseline) 683M 0 683M 2.73 GB global (ours) 686M 303M 989M 3.96 GB gate (ours) $\approx 0.42$M $\approx 1.67$B $\approx 1.67$B $\approx 6.7$ GB
Setup GPU(s) Peak memory (of 141 GB) Wall-clock time Data scale local (baseline) 1 $\times$ H200 15% ( $\approx$21 GB) $\sim$5 h 1.60M train / 0.45M val (16 classes) global (ours) 1 $\times$ H200 40% ( $\approx$56 GB) $\sim$22 h 1.60M train / 0.45M val (16 classes) gate (ours) 1 $\times$ H200 69% ( $\approx$97 GB) $\sim$5 h 1.60M train / 0.45M val (16 classes)
Counting policy for Table A6 (why Gate shows 1.67B non-trainable).
Table A6 reports parameter counts from PyTorch Lightning “summary” under the modules resident in memory during each stage. Gate training instantiates both experts end-to-end and freezes them: Path A uses UNI2-h, Path B uses UNI2-h and DINOv3-L/16 for context. Hence the non-trainable pool is
$\underbrace{\text{UNI2-h}}_{\text{Path local}}+\underbrace{\text{UNI2-h}}_{\text{Path global}}+\underbrace{\text{DINOv3-L/16}}_{\text{Path global}}\ \approx\ 2\times 683\text{M}+303\text{M}=1{,}669\text{M},$
matching the Gate row (non-trainable/total/size). Only the small Gate MLP and heads ($\sim 0.42$M) are trainable.
Appendix C Dataset composition and cohort splits
Cohort (organ) Cells Unique types Ovarian_wsi (ovary) 384,805 7 Pancreas_wsi (pancreas) 187,522 11 Heart-MultiCancerPanel (heart) 5,859 2 Kidney-MultiCancerPanel-Cancer (kidney) 27,550 5 Liver-MultiCancerPanel-Cancer (liver) 96,716 4 Lung-Cancer-v1+5k (lung) 195,133 7 Colon_Preview (colon) 446,933 5 Breast_Entire-Sample-Area-S1 (breast) 701,220 5 Combined 2,045,738 16
Table A6: Training cohorts after mapping/filtering.
Cohort (organ) Cells Unique types Pancreatic_Multimodal_Cell_Segmentation (pancreas) 74,689 6 Ovarian-ImmunoOnco+CustomAddon (ovary) 150,748 3 Lung-ImmunoOnco+CustomAddon (lung) 61,490 4 Combined 286,927 9
Table A7: Hold-out cohorts.
Class Top 10 marker genes Epithelial_Malignant EPCAM, KRT8, KRT18, KRT19, KRT17, KRT5, MKI67, MCM2, TOP2A, BIRC5 Cancer_Associated_Fibroblast COL1A1, COL1A2, COL3A1, ACTA2, TAGLN, PDGFRB, FAP, PDPN, POSTN, CXCL12 Macrophage CD68, LST1, CSF1R, MS4A7, C1QA, C1QB, C1QC, CD163, MRC1, ITGAM T_Cell CD3D, CD3E, CD2, CD4, CD8A, TRAC, CD247, LCK, IL7R, CCR7 Endothelial PECAM1, VWF, KDR, CD34, EMCN, FLT1, KLF2, CLDN5, ENG, ESAM Smooth_Muscle ACTA2, TAGLN, MYH11, CNN1, MYL9, DES, TPM2, CALD1, RGS5, MYLK Mast_Cell KIT, TPSAB1, TPSB2, CPA3, MS4A2, HDC, GATA2, FCER1A, SRGN, SLC18A2 Fibroblast COL1A1, COL1A2, DCN, LUM, PDGFRA, COL3A1, COL5A1, FBLN1, THY1, VIM Endothelial_CancerAssociated PECAM1, VWF, KDR, ICAM1, SELE, FLT1, VCAM1, ANGPT2, IL6, CXCL8 B_Cell MS4A1, CD79A, CD79B, CD74, CD19, CD22, CD37, CD40, BANK1, PAX5 Acinar PRSS1, PRSS3, CPA1, CPB1, CELA3A, CTRB1, CTRC, CEL, CPA2, REG1A Endocrine INS, GCG, SST, PPY, CHGA, IAPP, PCSK1, PCSK2, PAX6, MAFA Epithelial_Alveolar SFTPA1, SFTPA2, SFTPB, SFTPC, SLC34A2, AGER, LPCAT1, ABCA3, KRT8, KRT18 Epithelial_Airway KRT5, KRT8, KRT18, MUC1, SCGB1A1, KRT14, KRT15, KRT17, FOXJ1, EPCAM Ductal KRT7, KRT8, KRT18, KRT19, MUC1, EPCAM, KRT23, SOX9, CFTR, SPP1 epithelial EPCAM, KRT8, KRT18, KRT19, KRT5, KRT14, KRT15, KRT17, CDH1, KRT7 lymphocyte PTPRC, CD3D, CD3E, MS4A1, CD79A, CD4, CD8A, NKG7, GNLY, TRAC plasma MZB1, XBP1, SDC1, SLAMF7, PRDM1, IGKC, IGHG1, JCHAIN, TNFRSF17, CD38
Table A8: Class-level marker genes used for cell type annotation. We list the top 10 marker genes per class here for brevity; in our actual analysis we use the top 50 marker genes.
Appendix D Cohort-specific projection matrices
Definition and notation (unified).
Let $C_{\text{train}}$ and $C_{\text{eval}}$ be the numbers of classes in the training and evaluation taxonomies for a given cohort. We define a fixed binary projection
$M\in\{0,1\}^{C_{\text{train}}\times C_{\text{eval}}},\qquad M_{ij}=1\ \text{iff training class }i\ \text{maps to evaluation class }j.$
Given $p_{\text{train}}\in\Delta^{C_{\text{train}}-1}$, the evaluation-space probabilities are
$p_{\text{eval}}\;=\;M^{\top}\,p_{\text{train}}\;\in\;\mathbb{R}^{C_{\text{eval}}},$ (A1)
so $\big(p_{\text{eval}}\big)_{j}=\sum_{i:\,M_{ij}=1}\big(p_{\text{train}}\big)_{i}$. We rebuild labels so that each retained training class maps to exactly one evaluation class; thus $\sum_{j}M_{ij}\in\{0,1\}$ and $\sum_{j}(p_{\text{eval}})_{j}=1$.
Minimal reproducible example (shapes and a numeric toy).
Suppose $C_{\mathrm{train}}=5$ and $C_{\mathrm{eval}}=4$, where two epithelial subclasses are merged at evaluation:
$M\;=\;\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\in\{0,1\}^{5\times 4},\quad p_{\mathrm{train}}=\begin{bmatrix}0.10\\ 0.15\\ 0.35\\ 0.25\\ 0.15\end{bmatrix}\Rightarrow\ p_{\mathrm{eval}}=M^{\top}p_{\mathrm{train}}=\begin{bmatrix}0.10\\ 0.50\\ 0.25\\ 0.15\end{bmatrix}$.
Cohort matrices.
For each held-out cohort (lung, ovary, pancreas), we instantiate $M$ with shape $C_{\text{train}}\!\times\!C_{\text{eval}}$ exactly as printed below; rows enumerate training classes in the canonical taxonomy and columns enumerate that cohort’s evaluation classes (intersection set). The resulting $p_{\text{eval}}=M^{\top}p_{\text{train}}$ is used consistently in all tables and CI summaries.
Lung.
$M_{\text{lung}}\in\{0,1\}^{6\times 4}=\left[\begin{array}[]{c|cccc}&\text{Endothelial}&\text{Epithelial\_Malignant}&\text{T\_Cell}&\text{Tumor\_Associated\_Fibroblast}\\ \hline\cr\text{Epithelial\_Malignant}&0&1&0&0\\ \text{Endothelial}&1&0&0&0\\ \text{Fibroblast}&0&0&0&0\\ \text{Macrophage}&0&0&0&0\\ \text{T\_Cell}&0&0&1&0\\ \text{Tumor\_Associated\_Fibroblast}&0&0&0&1\\ \end{array}\right].$
Ovary.
$M_{\text{ovary}}\in\{0,1\}^{4\times 3}=\left[\begin{array}[]{c|ccc}&\text{Endothelial}&\text{Epithelial\_Malignant}&\text{Macrophage}\\ \hline\cr\text{Epithelial\_Malignant}&0&1&0\\ \text{Fibroblast}&0&0&0\\ \text{Macrophage}&0&0&1\\ \text{Endothelial}&1&0&0\\ \end{array}\right].$
Pancreas.
$M_{\text{pancreas}}\in\{0,1\}^{7\times 6}=\left[\begin{array}[]{c|cccccc}&\text{Acinar}&\text{Ductal}&\text{Endocrine}&\text{Endothelial}&\text{Fibroblast\_like}&\text{T\_Cell}\\ \hline\cr\text{Ductal}&0&1&0&0&0&0\\ \text{Fibroblast}&0&0&0&0&1&0\\ \text{Acinar}&1&0&0&0&0&0\\ \text{Tumor\_Associated\_Fibroblast}&0&0&0&0&1&0\\ \text{Endothelial}&0&0&0&1&0&0\\ \text{T\_Cell}&0&0&0&0&0&1\\ \text{Endocrine}&0&0&1&0&0&0\\ \end{array}\right].$
Appendix E Cohort-wise Macro/Micro-F1 and Accuracy with 95% Confidence Intervals
We report macro-F1 and micro-F1 for all cohorts using non-parametric bootstrap (B=1000B{=}1000 resamples per cohort). All entries are formatted as mean (lower, upper), where values in parentheses denote the 95% confidence interval; best in each row is in bold. Evaluation strictly follows $p_{\text{eval}}=M^{\top}p_{\text{train}}$.
Table A9: Macro-F1 (mean (lower, upper), 95% CI).
Cohort DINO-B/16 DINO-L/16 DINO-S+/16 LOKI PLIP MUSK Path Local Path Global Fusion Gate Lung 0.4500 (0.4465, 0.4535) 0.4316 (0.4281, 0.4351) 0.2663 (0.2628, 0.2698) 0.4839 (0.4804, 0.4874) 0.3945 (0.3910, 0.3980) 0.5290 (0.5255, 0.5325) 0.5930 (0.5883, 0.5981) 0.6658 (0.6610, 0.6684) 0.6971 (0.6913, 0.7026) Ovary 0.4051 (0.4016, 0.4086) 0.4056 (0.4021, 0.4091) 0.3590 (0.3555, 0.3625) 0.3696 (0.3661, 0.3731) 0.4652 (0.4617, 0.4687) 0.5567 (0.5532, 0.5602) 0.6337 (0.6296, 0.6377) 0.6743 (0.6728, 0.6762) 0.6910 (0.6867, 0.6948) Pancreas 0.1167 (0.1132, 0.1202) 0.1033 (0.0998, 0.1068) 0.1667 (0.1632, 0.1702) 0.1337 (0.1302, 0.1372) 0.0763 (0.0728, 0.0798) 0.2308 (0.2273, 0.2343) 0.5201 (0.5170, 0.5232) 0.4801 (0.4790, 0.4808) 0.4811 (0.4772, 0.4847) Overall 0.3239 (0.3204, 0.3274) 0.3135 (0.3100, 0.3170) 0.2640 (0.2605, 0.2675) 0.3291 (0.3256, 0.3326) 0.3120 (0.3085, 0.3155) 0.4400 (0.4365, 0.4435) 0.5300 (0.5275, 0.5326) 0.5367 (0.5364, 0.5371) 0.6235 (0.6191, 0.6276)
Table A10: Micro-F1 (mean (lower, upper), 95% CI). For single-label classification, Micro-F1 equals accuracy.
Cohort DINO-B/16 DINO-L/16 DINO-S+/16 LOKI PLIP MUSK Path Local Path Global Fusion Gate Lung 0.6651 (0.6613, 0.6689) 0.6413 (0.6376, 0.6451) 0.3166 (0.3129, 0.3203) 0.7013 (0.6974, 0.7049) 0.4883 (0.4845, 0.4923) 0.6728 (0.6688, 0.6763) 0.8494 (0.8466, 0.8521) 0.8611 (0.8594, 0.8622) 0.8695 (0.8669, 0.8721) Ovary 0.8153 (0.8134, 0.8173) 0.8213 (0.8193, 0.8233) 0.7048 (0.7025, 0.7071) 0.4973 (0.4948, 0.4998) 0.7713 (0.7692, 0.7735) 0.7266 (0.7244, 0.7290) 0.8910 (0.8894, 0.8924) 0.7452 (0.7446, 0.7463) 0.8932 (0.8916, 0.8949) Pancreas 0.2451 (0.2427, 0.2478) 0.2521 (0.2495, 0.2546) 0.2131 (0.2105, 0.2157) 0.2199 (0.2174, 0.2223) 0.1587 (0.1566, 0.1609) 0.2615 (0.2589, 0.2643) 0.4967 (0.4937, 0.4998) 0.4722 (0.4703, 0.4733) 0.5003 (0.4973, 0.5036) Overall 0.5995 (0.5979, 0.6012) 0.6001 (0.5985, 0.6018) 0.4683 (0.4665, 0.4699) 0.4464 (0.4445, 0.4482) 0.5158 (0.5141, 0.5177) 0.5640 (0.5623, 0.5658) 0.7539 (0.7525, 0.7554) 0.7452 (0.7446, 0.7463) 0.7601 (0.7578, 0.7625)
E.1 Does local-aware Gating Benefit the Global Path?
Evaluation Setup
We evaluate the effect of the local-aware gating signal specifically on the Global Path. All results are computed on the 16-class intersection.
Finding
As shown in Table A11, local-aware consistently improves Path global across all metrics:
  • Overall accuracy improves by +0.013+0.013 and macro F1 by +0.022+0.022.
  • Calibration is noticeably better: ECE decreases by $0.013$ and Brier score decreases by $0.019$.
  • On challenging subsets, macro F1 increases by +0.026+0.026 (low-confidence) and +0.032+0.032 (high-entropy).
  • AUROC also shows a slight positive change (+0.002+0.002).
These results indicate that local-aware gating provides a beneficial confidence prior that stabilizes and strengthens the Global Path’s predictions, even though the mechanism originates from the path local.
[…]
local_aware_ON 0.856 0.773 0.045 0.210 0.990 0.500 0.620
local_aware_OFF 0.842 0.751 0.058 0.229 0.988 0.475 0.589
$\Delta$ (ON–OFF) +0.013 +0.022 -0.013 -0.019 +0.002 +0.026 +0.032
Additional visualizations
Figure A1: Additional Grad-CAM visualizations (Path-local expert).
Figure A2: Additional Grad-CAM visualizations for LOKI, MUSK, and PLIP (Ovary).
Acknowledgements
As future work, we plan to evaluate zero-/few-shot transfer on non-Xenium, public H&E datasets with coarse label mappings, to further assess cross-panel, cross-stain, and cross-institution generalization. We emphasize that the present study already establishes strong evidence: Xenium provides subcellular-resolution, high-plex spatial gene expression with cell-level coordinates, enabling objective, reproducible labels that serve as a high-quality supervision signal for cell-wise modeling. Our results demonstrate that, under this gold-standard molecular supervision, NuClass succeeds at subcellular granularity and yields calibrated, interpretable predictions across fully held-out cohorts.